OpenStudy (ayonnaleflore):

Find the zeros of the polynomial function. f(x) = x2 – 12x + 20 4, 5 2, 10 –4, –5 –2, –10

OpenStudy (wcrmelissa2001):

by zeros do you mean roots?

OpenStudy (ayonnaleflore):

now this one i got totally lost on

OpenStudy (ayonnaleflore):

yes

OpenStudy (wcrmelissa2001):

https://www.mathsisfun.com/algebra/factoring-quadratics.html Sorry, i'm not sure how to explain but this is basically how it works. LIke er... the link should be more helpful... @mathmale

OpenStudy (mathmale):

f(x) = x2 – 12x + 20 Should be written as f(x) = x^2 – 12x + 20. What are some of the methods of factoring quadratics that you've learned and used before? 1. factoring by grouping 2. quadratic formula 3. completing the square 4. graphing 5. others? Why not try some of these?

OpenStudy (ayonnaleflore):

so how would i do it if i wanted to use the Quadratic formula

OpenStudy (mathmale):

Identify the coefficients a, b qnd c of f(x) = x^2 – 12x + 20. a= b= c= Plug these values into the quadratic formula\[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] Show all work and share your results, please.

OpenStudy (1sttime4everything):

mathmale can i ask you something

OpenStudy (mathmale):

What would you like to ask?

OpenStudy (ayonnaleflore):

a=1 b=-12 c=20

OpenStudy (mathmale):

Nice work. x=??

OpenStudy (mathmale):

What result have you obtained?